# Generalized Moving Least-Squares for Solving Vector-valued PDEs on Unknown Manifolds

### v1

@@ 6 Conclusions and discussion
In this paper, we extended the GMLS approach using both intrinsic formulation and extrinsic formulation for approximating vector Laplacians and covariant derivative and further for solving vector-valued PDEs on unknown smooth 2D manifolds, identified by a randomly-sampled data point cloud. The scheme can reach an arbitrary-order algebraic accuracy on smooth manifolds with small curvatures. For the intrinsic method, we approximated the geometry quantities, functions, and their derivatives in the local Monge coordinate system to simplify the intrinsic formulation of the Bochner Laplacian in (3.2.2). Then we transformed the representation of GMLS weights from the local tangent-vector basis to the global tangent-vector basis in order to assemble these weights into the sparse $2N\times 2N$ Bochner Laplacian matrix. For the extrinsic method, we saved the memory cost and computational cost in approximating the Bochner Laplacian and covariant derivative, which improved our previous results in [22]. In particular, we applied the GMLS approach to make the discretization sparse and we also reduced the dimension of Laplacian matrices as well as the nonzero entries in the matrices by applying an appropriate projection.
There are several remaining questions for future investigation. First, the stability problem still exists in the GMLS method for approximating the vector Laplacians on manifolds with high curvatures (such as bumpy spheres discussed in Section 5.3). Second, the GMLS approach whose hypothesis space only contains multivariate polynomials may fail to approximate discontinuous solutions such as shock waves of nonlinear Burgers’ equation. Third, it is of interest to solve vector-valued PDEs with classical boundary conditions on manifolds with boundaries identified with randomly sampled data points.
@@ Acknowledgments
S. J. was supported by the NSFC Grant No. 12101408. Q. Y. was supported by the Simons Foundation grant 601937, DNA. We would like to thank John Harlim for helpful comments.
@@ Appendix A Intrinsic formulation for other vector Laplacians
In this section, we first review some concepts and notations from intrinsic differential geometry [30, 55]. Then we discuss the intrinsic GMLS approximation of the L Laplacian (which involves the symmetric stress tensor) and Hodge Laplacian (which involves the anti-symmetric stress tensor) acting on vector fields.
@@ A.1 Intrinsic Differential Geometry
Let $M$ be a $d$-dimensional Riemannian manifold with a local coordinate $(\xi^{1},..,\xi^{d})$ defined on an open subset $O$ in $M$. The Riemannian metric is $g=g_{ij}d\xi^{i}\otimes d\xi^{j}$ with the components $g_{ij}=\langle\partial_{i},\partial_{j}\rangle,$ where $\partial_{i}=\frac{\partial}{\partial\xi^{i}}$ is the $i$th coordinate vector field and $(d\xi^{1},...,d\xi^{d})$ is the associated coordinate coframe. Recall that $d\xi^{j}(\partial_{i})=\delta_{i}^{j}$, where $\delta_{i}^{j}$ is the delta function.
To define the vector Laplacian, we first introduce the concepts for the gradient operator and the divergence operator. Let $u=\tilde{u}^{i}\partial_{i}\in\mathfrak{X}(M)$ be a vector field and let $\nabla$ be the Levi-Civita connection. Then the covariant differentiation is defined as
$\nabla u=\tilde{u}_{;j}^{i}\frac{\partial}{\partial\xi^{i}}\otimes d\xi^{j}$
@@
${\mathrm{grad}}_{g}u={\sharp\nabla u}=g^{kj}\tilde{u}_{;k}^{i}\frac{\partial}{\partial\xi^{i}}\otimes\frac{\partial}{\partial\xi^{j}}.$
For a vector field $u$, its divergence can be expressed as
$\mathrm{div}_{g}\left(u\right)=\text{tr}\left(\nabla u\right)=\tilde{u}_{;i}^{i}.$
Let $v=\tilde{v}^{ij}\partial_{i}\otimes\partial_{j}\in\mathfrak{X}(M)\times\mathfrak{X}(M)$ be a $(2,0)$ tensor field. One can define the divergence of a $(2,0)$ tensor field (see e.g. Ch 2 in [55]):
$\mathrm{div}_{g}(v):=\mathrm{tr}_{1}^{2}\left(\nabla v\right)=\mathrm{tr}_{1}^{2}\left(\tilde{v}_{;k}^{ij}\frac{\partial}{\partial\xi^{i}}\otimes\frac{\partial}{\partial\xi^{j}}\otimes d\xi^{k}\right)=\tilde{v}_{;k}^{ik}\partial_{i},$ (42)
where the trace operator $\mathrm{tr}^{2}_{1}$ is taken on the last two indices of the tensor field $\nabla v$.
@@
1) The Bochner Laplacian acting on a vector field $u=\tilde{u}^{k}\partial_{k}$ is defined as
$\Delta_{B}u:=\mathrm{div}_{g}({\mathrm{grad}}_{g}u)=\mathrm{div}_{g}\left(g^{kj}\tilde{u}_{;k}^{i}\frac{\partial}{\partial\xi^{i}}\otimes\frac{\partial}{\partial\xi^{j}}\right)=g^{kj}\tilde{u}_{;kj}^{i}\partial_{i}.$ (43)
@@
$S:=\frac{1}{2}\left({\mathrm{grad}}_{g}u+({\mathrm{grad}}_{g}u)^{\top}\right)=\frac{1}{2}\left(g^{mj}\tilde{u}_{;m}^{i}+g^{mi}\tilde{u}_{;m}^{j}\right)\partial_{i}\otimes\partial_{j}.$
Then the (negative-definite) L Laplacian acting on the vector field $u$ can be defined and computed in the local coordinates as
$\Delta_{L}u:=2\mathrm{div}_{g}\left(S\right)=\mathrm{div}_{g}\left({\mathrm{grad}}_{g}u+({\mathrm{grad}}_{g}u)^{\top}\right)=\left(g^{ij}\tilde{u}_{;ij}^{k}+g^{ki}\tilde{u}_{;ij}^{j}\right)\partial_{k}.$ (44)
3) The (negative-definite) Hodge Laplacian acting on the vector field $u=\tilde{u}^{k}\partial_{k}$ is defined as
$\Delta_{H}u:=-\sharp(\mathrm{d}^{\ast}\mathrm{d}+\mathrm{d}\mathrm{d}^{\ast})\flat u,$
@@
$\Delta_{H}u$ $=$ $\mathrm{div}_{g}({\mathrm{grad}}_{g}u-({\mathrm{grad}}_{g}u)^{\top})+{\mathrm{grad}}_{g}\left(\mathrm{div}_{g}\left({u}\right)\right)$ (45)
@@ A.2 Intrinsic GMLS approximation of other vector Laplacians
In this section, we first provide the representation of $\Delta_{L}u$ and $\Delta_{H}u$ in the Monge coordinate system following the similar computational procedure for the Bochner $\Delta_{B}u$ in Section 3.2.1. Recall that the notations, $\{a_{\alpha_{1},\alpha_{2}}|2\leq\alpha_{1}+\alpha_{2}\leq l\}$, are the regression coefficients of the parameterized surface (see equation (9)), and the notations, $\beta_{\alpha_{1},\alpha_{2}}$ and $\gamma_{\alpha_{1},\alpha_{2}}$ for $0\leq\alpha_{1}+\alpha_{2}\leq l$, are the coefficients of respective component, $\tilde{u}^{1}$ and $\tilde{u}^{2}$, of the vector field $u=\tilde{u}^{k}\partial_{k}$ (see equation (17)).
Then the L Laplacian in (44) at the base point $\mathbf{x}_{0}$ can be approximated as
$\Delta_{L}u(\mathbf{x}_{0})$
@@
$+\left(\beta_{1,1}+\beta_{0,0}[4a_{2,0}a_{1,1}+4a_{0,2}a_{1,1}]+4\gamma_{0,2}+2\gamma_{2,0}+\gamma_{0,0}\left[8(a_{0,2})^{2}+(a_{1,1})^{2}+4a_{2,0}a_{2,0}\right]\right)\partial_{2}|_{\mathbf{x}_{0}}.$
The Hodge Laplacian in (45) can be approximated as
$\Delta_{H}u(\mathbf{x}_{0})=(g^{ij}\tilde{u}_{;ij}^{k}-g^{kj}\tilde{u}_{;ji}^{i}+g^{jk}\tilde{u}_{;ij}^{i})\partial_{k}|_{\mathbf{x}_{0}}$
@@
Then one can follow the similar GMLS procedure in Section 3.2 to approximate $\Delta_{L}u$ and $\Delta_{H}u$ on a point cloud.
@@ B.1 Bochner Laplacian
Here, we supplement a detailed derivation of the extrinsic formulation for the Bochner Laplacian in Section 4.1. Let $v=v^{jk}\frac{\partial}{\partial\xi^{j}}\otimes\frac{\partial}{\partial\xi^{k}}\in\mathfrak{X}(M)\times\mathfrak{X}(M)$ be a $(2,0)$ tensor field, and let
$V=\frac{\partial X^{s}}{\partial\xi^{j}}v^{jk}\frac{\partial X^{t}}{\partial\xi^{k}}\frac{\partial}{\partial X^{s}}\otimes\frac{\partial}{\partial X^{t}}:=V^{st}\frac{\partial}{\partial X^{s}}\otimes\frac{\partial}{\partial X^{t}},$
be a smooth extension of $v$ onto an open subset in $\mathbb{R}^{3}$ such that $V|_{M}=v$. In general, the tensor field $v$ is not symmetric, that is, $v^{jk}\neq v^{kj}$. We now consider the extrinsic formulation for the divergence of a $(2,0)$ tensor field $v$. Using the definition in (42), one can verify the identity extension formulation for the divergence (see e.g. [22]),
$\mathrm{div}_{g}\left(v\right)=\mathbf{P}\mathrm{tr}^{2}_{1}\left(\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}\left(V\right)\right)=\mathbf{P}\mathrm{tr}^{2}_{1}\left(\mathcal{G}_{r}V^{st}\frac{\partial}{\partial X^{s}}\otimes\frac{\partial}{\partial X^{t}}\otimes\mathrm{d}X^{r}\right)=\mathbf{P}\left(\mathcal{G}_{t}V^{st}\frac{\partial}{\partial X^{s}}\right),$
@@
$\mathrm{div}_{g}\left(v\right)$ $=$ $\mathbf{P}\left[\begin{array}[]{c}\sum_{k=1}^{3}\mathcal{G}_{k}V^{1k}\\ \sum_{k=1}^{3}\mathcal{G}_{k}V^{2k}\\ \sum_{k=1}^{3}\mathcal{G}_{k}V^{3k}\end{array}\right]=\mathbf{P}\left[\begin{array}[]{ccc}\mathcal{G}_{1}&&\\ &\mathcal{G}_{1}&\\ &&\mathcal{G}_{1}\end{array}\right]\left[\begin{array}[]{c}V^{11}\\ V^{21}\\ V^{31}\end{array}\right]+\cdots+\mathbf{P}\left[\begin{array}[]{ccc}\mathcal{G}_{3}&&\\ &\mathcal{G}_{3}&\\ &&\mathcal{G}_{3}\end{array}\right]\left[\begin{array}[]{c}V^{13}\\ V^{23}\\ V^{33}\end{array}\right]$ (61)
$=$ $\mathcal{H}_{1}\left[\begin{array}[]{c}V^{11}\\ V^{21}\\ V^{31}\end{array}\right]+\mathcal{H}_{2}\left[\begin{array}[]{c}V^{12}\\ V^{22}\\ V^{32}\end{array}\right]+\mathcal{H}_{3}\left[\begin{array}[]{c}V^{13}\\ V^{23}\\ V^{33}\end{array}\right]:=\mathcal{H}_{1}V\mathbf{|}_{1}+\mathcal{H}_{2}V\mathbf{|}_{2}+\mathcal{H}_{3}V\mathbf{|}_{3}$ (71)
where $\mathcal{H}_{s}$ is defined as in (25) and $V\mathbf{|}_{s}:=(V^{1s},V^{2s},V^{3s})^{\top}$ is the $s$th column of tensor field $V$ for $s=1,2,3$.
Let $v={\mathrm{grad}}_{g}u\in\mathfrak{X}(M)\times\mathfrak{X}(M)$ be the gradient of a vector field which is a $(2,0)$ tensor field. Using the extrinsic formulation in (4.1), one obtains the equation:
$V=\left(V\mathbf{|}_{1},V\mathbf{|}_{2},V\mathbf{|}_{3}\right)=\left(\mathcal{H}_{1}U,\mathcal{H}_{2}U,\mathcal{H}_{3}U\right)=\mathbf{P}\left(\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}U\right)\mathbf{P},$
@@
$V\mathbf{|}_{k}=\mathcal{H}_{k}U,\text{ \ \ for }k=1,2,3\text{.}$ (72)
Combining (71) and (72), one can obtain the extrinsic formulation for the Bochner Laplacian of a vector field, $\Delta_{B}u=g^{ij}u_{;ij}^{k}\frac{\partial}{\partial\xi^{k}}=\mathrm{div}_{g}({\mathrm{grad}}_{g}{u)}$, as
$\Delta_{B}u$ $=$ $\mathrm{div}_{g}({\mathrm{grad}}_{g}u)=\mathrm{div}_{g}\left(v\right)=\mathcal{H}_{1}V\mathbf{|}_{1}+\mathcal{H}_{2}V\mathbf{|}_{2}+\mathcal{H}_{3}V\mathbf{|}_{3}$
$=$ $\mathcal{H}_{1}\mathcal{H}_{1}U+\mathcal{H}_{2}\mathcal{H}_{2}U+\mathcal{H}_{3}\mathcal{H}_{3}U:=\bar{\Delta}_{B}U.$

### v2

@@ 6 Conclusions and discussion
In this paper, we extended the GMLS approach using both intrinsic formulation and extrinsic formulation for approximating vector Laplacians and covariant derivative and further for solving vector-valued PDEs on unknown smooth manifolds, identified by a randomly-sampled data point cloud. The scheme can reach an arbitrary-order algebraic accuracy on smooth manifolds. For the intrinsic method, we approximated the geometric quantities, functions, and their derivatives in the local Monge coordinate system to simplify the intrinsic formulation of the Bochner Laplacian in (3.2.2) and (25). Then we transformed the representation of GMLS weights from the local tangent-vector basis to the global tangent-vector basis in order to assemble these weights into the sparse $dN\times dN$ Bochner Laplacian matrix. For the extrinsic method, we saved the memory cost and computational cost in approximating the Bochner Laplacian and covariant derivative, which improved our previous results in [24]. In particular, we applied the GMLS approach as a sparse discretization and we also reduced the dimension of Laplacian matrices as well as the nonzero entries in the matrices by applying an appropriate coordinate transformation.
Several open questions remain for future investigation. For our GMLS method, the total complexity is $O(Nn)$, allowing it to accommodate relatively large ambient dimensions. However, the requirements for storage, stability and convergence impose a strong restriction on the intrinsic dimension $d$. For $d\leq 3$, as discussed in the 3D example in subsection 5.2.3, an interpolation approach, in particular, Radial Basis Function-Finite Difference (RBF-FD) [15, 29], could potentially improve the stability and convergence. Since the approximation space of the RBF-FD approach contains both multivariate polynomials and radial basis functions, we expect it to outperform the GMLS approach, whose approximation space consists solely of multivariate polynomials. For even larger intrinsic dimensions ($d\geq 4$), both GMLS and RBF-FD approaches may not be efficient for solving vector-valued PDE problems on manifolds due to the increased data requirements and memory storage demands. In such cases, neural networks could be considered for discretization and approximation of PDE solutions, while the framework for approximating vector Laplacians developed in this work is still useful. Additionally, an important direction for future work is solving vector-valued PDEs with classical boundary conditions on manifolds with boundaries.
@@ Acknowledgments
S. J. was supported by the NSFC Grant No. 12471412 and the HPC Platform of ShanghaiTech University. Q. Y. was supported by the Simons Foundation grant 601937, DNA. We would like to thank John Harlim for helpful comments.
@@ Appendix A Intrinsic formulation for other vector Laplacians
In this section, we first review some concepts and notations from intrinsic differential geometry [33, 60]. Then we discuss the intrinsic GMLS approximation of the L Laplacian (which involves the symmetric stress tensor) and the Hodge Laplacian (which involves the anti-symmetric stress tensor) acting on vector fields. In addition, we also present the proof for Proposition 3.2.
@@ A.1 Intrinsic Differential Geometry
Let $M$ be a $d$-dimensional Riemannian manifold with a local coordinate $(\xi^{1},..,\xi^{d})$ defined on an open subset $O$ in $M$. The Riemannian metric is $\bm{g}=g_{ij}d\xi^{i}\otimes d\xi^{j}$ with the components $g_{ij}=\langle\partial_{i},\partial_{j}\rangle,$ where $\partial_{i}=\frac{\partial}{\partial\xi^{i}}$ is the $i$th coordinate vector field and $(d\xi^{1},...,d\xi^{d})$ is the associated coordinate coframe. Recall that $d\xi^{j}(\partial_{i})=\delta_{i}^{j}$, where $\delta_{i}^{j}$ is the delta function.
To define the vector Laplacian, we first introduce the concepts for the gradient operator and the divergence operator. Let $\bm{u}=\tilde{u}^{i}\partial_{i}\in\mathfrak{X}(M)$ be a vector field and let $\nabla$ be the Levi-Civita connection. Then the covariant differentiation is defined as
$\nabla\bm{u}=\tilde{u}_{;j}^{i}\frac{\partial}{\partial\xi^{i}}\otimes d\xi^{j}$
@@
${\mathrm{grad}}_{g}\bm{u}={\sharp\nabla\bm{u}}=g^{kj}\tilde{u}_{;k}^{i}\frac{\partial}{\partial\xi^{i}}\otimes\frac{\partial}{\partial\xi^{j}}.$
For a vector field $\bm{u}$, its divergence can be expressed as
$\mathrm{div}_{g}\left(\bm{u}\right)=\text{tr}\left(\nabla\bm{u}\right)=\tilde{u}_{;i}^{i}.$
Let $\bm{v}=\tilde{v}^{ij}\partial_{i}\otimes\partial_{j}\in\mathfrak{X}(M)\times\mathfrak{X}(M)$ be a $(2,0)$ tensor field. One can define the divergence of a $(2,0)$ tensor field (see e.g. Ch 2 in [60]):
$\mathrm{div}_{g}(\bm{v}):=\mathrm{tr}_{1}^{2}\left(\nabla\bm{v}\right)=\mathrm{tr}_{1}^{2}\left(\tilde{v}_{;k}^{ij}\frac{\partial}{\partial\xi^{i}}\otimes\frac{\partial}{\partial\xi^{j}}\otimes d\xi^{k}\right)=\tilde{v}_{;k}^{ik}\partial_{i},$ (56)
where the trace operator $\mathrm{tr}^{2}_{1}$ is taken on the last two indices of the tensor field $\nabla\bm{v}$.
@@
1) The Bochner Laplacian acting on a vector field $\bm{u}=\tilde{u}^{k}\partial_{k}$ is defined as
$\Delta_{B}\bm{u}:=\mathrm{div}_{g}({\mathrm{grad}}_{g}\bm{u})=\mathrm{div}_{g}\left(g^{kj}\tilde{u}_{;k}^{i}\frac{\partial}{\partial\xi^{i}}\otimes\frac{\partial}{\partial\xi^{j}}\right)=g^{kj}\tilde{u}_{;kj}^{i}\partial_{i}.$ (57)
@@
$S:=\frac{1}{2}\left({\mathrm{grad}}_{g}\bm{u}+({\mathrm{grad}}_{g}\bm{u})^{\top}\right)=\frac{1}{2}\left(g^{mj}\tilde{u}_{;m}^{i}+g^{mi}\tilde{u}_{;m}^{j}\right)\partial_{i}\otimes\partial_{j}.$
Then the (negative-definite) L Laplacian acting on the vector field $\bm{u}$ can be defined and computed in the local coordinates as
$\Delta_{L}\bm{u}:=2\mathrm{div}_{g}\left(S\right)=\mathrm{div}_{g}\left({\mathrm{grad}}_{g}\bm{u}+({\mathrm{grad}}_{g}\bm{u})^{\top}\right)=\left(g^{ij}\tilde{u}_{;ij}^{k}+g^{ki}\tilde{u}_{;ij}^{j}\right)\partial_{k}.$ (58)
3) The (negative-definite) Hodge Laplacian acting on the vector field $\bm{u}=\tilde{u}^{k}\partial_{k}$ is defined as
$\Delta_{H}\bm{u}:=-\sharp(\mathrm{d}^{\ast}\mathrm{d}+\mathrm{d}\mathrm{d}^{\ast})\flat\bm{u},$
@@
$\Delta_{H}\bm{u}$ $=$ $\mathrm{div}_{g}({\mathrm{grad}}_{g}\bm{u}-({\mathrm{grad}}_{g}\bm{u})^{\top})+{\mathrm{grad}}_{g}\left(\mathrm{div}_{g}\left(\bm{u}\right)\right)$ (59)
@@ A.2 Intrinsic GMLS approximation of other vector Laplacians
In this section, we first provide the representation of $\Delta_{L}\bm{u}$ and $\Delta_{H}\bm{u}$ in the Monge coordinate system following the similar computational procedure for the Bochner $\Delta_{B}\bm{u}$ in Section 3.2.1. Recall that the notations, $\{a_{\alpha_{1},\alpha_{2}}|2\leq\alpha_{1}+\alpha_{2}\leq l\}$, are the regression coefficients of the parameterized surface (see equation (10)), and the notations, $\beta_{\alpha_{1},\alpha_{2}}$ and $\gamma_{\alpha_{1},\alpha_{2}}$ for $0\leq\alpha_{1}+\alpha_{2}\leq l$, are the coefficients of respective component, $\tilde{u}^{1}$ and $\tilde{u}^{2}$, of the vector field $\bm{u}=\tilde{u}^{k}\partial_{k}$ (see equation (18)).
Then for $d=2$, the L Laplacian in (58) at the base point $\mathbf{x}_{0}$ can be approximated as
$\Delta_{L}\bm{u}(\mathbf{x}_{0})$
@@
$+\left(\beta_{1,1}+\beta_{0,0}[4a_{2,0}a_{1,1}+4a_{0,2}a_{1,1}]+4\gamma_{0,2}+2\gamma_{2,0}+\gamma_{0,0}\left[8(a_{0,2})^{2}+(a_{1,1})^{2}+4a_{2,0}a_{2,0}\right]\right)\partial_{2}|_{\mathbf{x}_{0}},$
and the Hodge Laplacian in (59) can be approximated as
$\Delta_{H}\bm{u}(\mathbf{x}_{0})=(g^{ij}\tilde{u}_{;ij}^{k}-g^{kj}\tilde{u}_{;ji}^{i}+g^{jk}\tilde{u}_{;ij}^{i})\partial_{k}|_{\mathbf{x}_{0}}$
@@
For the general $d$ dimensional manifold, one can obtain the L Laplacian
$\Delta_{L}\bm{u}(\mathbf{x}_{0})=\sum_{i=1}^{d}\left(\partial_{ii}\tilde{u}^{k}+\partial_{ki}\tilde{u}^{i}+\sum_{s=1}^{n-d}\tilde{u}^{l}\partial_{il}q^{s}\partial_{ki}q^{s}+\sum_{s=1}^{n-d}\tilde{u}^{l}\partial_{kl}q^{s}\partial_{ii}q^{s}\right)\partial_{k}|_{\mathbf{x}_{0}},$
and the Hodge Laplacian
$\Delta_{H}\bm{u}(\mathbf{x}_{0})=\sum_{i=1}^{d}\left(\partial_{ii}\tilde{u}^{k}+2\sum_{s=1}^{n-d}\tilde{u}^{l}\partial_{ik}q^{s}\partial_{il}q^{s}-\sum_{s=1}^{n-d}\tilde{u}^{l}\partial_{ii}q^{s}\partial_{kl}q^{s}\right)\partial_{k}|_{\mathbf{x}_{0}}.$
Then one can follow the similar GMLS procedure in Sections 3.2 and 3.3 to approximate $\Delta_{L}\bm{u}$ and $\Delta_{H}\bm{u}$ on a point cloud.
A.3 Proof for Proposition 3.2
Here we prove Proposition 3.2.
Proof.
From equation (27), we have
${\mathbf{x}}-{\mathbf{x}_{0}}=\sum_{s=1}^{n}p^{s}\mathbf{e}_{s},$
where $\left\{\mathbf{e}_{s}\right\}_{s=1}^{n}$ are the standard orthonormal bases of the Euclidean space. By comparing with the Monge coordinate (22), we arrive at
$\sum_{s=1}^{n}p^{s}\mathbf{e}_{s}=\sum_{i=1}^{d}\theta_{i}\bm{t}_{i}({\mathbf{x}_{0}})+\sum_{s=1}^{n-d}q^{s}\bm{n}_{s}({\mathbf{x}_{0}}),$
which implies that
$p^{s}=x^{s}-x_{0}^{s}=\sum_{i=1}^{d}\theta_{i}t_{i}^{s}+\sum_{j=1}^{n-d}q^{j}n_{j}^{s}.$ (60)
Here $\left\{t_{i}^{s}\right\}_{i=1,\ldots,d}^{s=1,\ldots,n}$ is the $s$th component of the $i$th tangent vector $\bm{t}_{i}$ and $\{n_{j}^{s}\}_{j=1,\ldots,n-d}^{s=1,\ldots,n}$ is the $s$th component of the $j$th normal vector $\bm{n}_{j}$  in the Monge coordinate system. Then we conclude that
$\sum_{s=1}^{n}\partial_{ir}p^{s}\partial_{kl}p^{s}$ $=$ $\sum_{s=1}^{n}\partial_{ir}\left(\sum_{v=1}^{d}\theta_{v}t_{v}^{s}+\sum_{j=1}^{n-d}q^{j}n_{j}^{s}\right)\partial_{kl}\left(\sum_{h=1}^{d}\theta_{h}t_{h}^{s}+\sum_{m=1}^{n-d}q^{m}n_{m}^{s}\right)$
$=$ $\sum_{s=1}^{n}\sum_{j=1}^{n-d}\sum_{m=1}^{n-d}(\partial_{ir}q^{j})n_{j}^{s}(\partial_{kl}q^{m})n_{m}^{s}=\sum_{j=1}^{n-d}\sum_{m=1}^{n-d}(\partial_{ir}q^{j})(\partial_{kl}q^{m})\sum_{s=1}^{n}n_{j}^{s}n_{m}^{s}$
$=$ $\sum_{j=1}^{n-d}\sum_{m=1}^{n-d}(\partial_{ir}q^{j})(\partial_{kl}q^{m})\delta_{jm}=\sum_{s=1}^{n-d}\partial_{ir}q^{s}\partial_{kl}q^{s}.$
Using the expresion (27) of $p^{s}$ and the quadratic form (28), we arrive at
$\partial_{ir}p^{s}|_{{\mathbf{x}_{0}}}=2c_{ir}^{s},\text{ \ \ for all }1\leq i,r\leq d\text{ and }1\leq s\leq n.$ (61)
Combining (24), (30) and (61), we obtain that
$\frac{\partial\Gamma_{kl}^{i}}{\partial\theta_{r}}|_{\mathbf{x}_{0}}=\sum_{s=1}^{n-d}\partial_{ir}q^{s}\partial_{kl}q^{s}|_{\mathbf{x}_{0}}=\sum_{s=1}^{n}\partial_{ir}p^{s}\partial_{kl}p^{s}|_{\mathbf{x}_{0}}=\sum_{s=1}^{n}4c_{ir}^{s}c_{kl}^{s}.$
@@ B.1 Bochner Laplacian
Here, we supplement a detailed derivation of the extrinsic formulation for the Bochner Laplacian in Section 4.1. Let $\bm{v}=v^{jk}\frac{\partial}{\partial\xi^{j}}\otimes\frac{\partial}{\partial\xi^{k}}\in\mathfrak{X}(M)\times\mathfrak{X}(M)$ be a $(2,0)$ tensor field, and let
$\bm{V}=\frac{\partial X^{s}}{\partial\xi^{j}}v^{jk}\frac{\partial X^{t}}{\partial\xi^{k}}\frac{\partial}{\partial X^{s}}\otimes\frac{\partial}{\partial X^{t}}:=V^{st}\frac{\partial}{\partial X^{s}}\otimes\frac{\partial}{\partial X^{t}},$
be a smooth extension of $\bm{v}$ onto an open subset in $\mathbb{R}^{3}$ such that $\bm{V}|_{M}=\bm{v}$. In general, the tensor field $\bm{v}$ is not symmetric, that is, $v^{jk}\neq v^{kj}$. We now consider the extrinsic formulation for the divergence of a $(2,0)$ tensor field $\bm{v}$. Using the definition in (56), one can verify the identity extension formulation for the divergence (see e.g. [24]),
$\mathrm{div}_{g}\left(\bm{v}\right)=\mathbf{P}\mathrm{tr}^{2}_{1}\left(\mathbf{P}\bar{\nabla}_{\mathbb{R}^{3}}\left(\bm{V}\right)\right)=\mathbf{P}\mathrm{tr}^{2}_{1}\left(\mathcal{G}_{r}V^{st}\frac{\partial}{\partial X^{s}}\otimes\frac{\partial}{\partial X^{t}}\otimes\mathrm{d}X^{r}\right)=\mathbf{P}\left(\mathcal{G}_{t}V^{st}\frac{\partial}{\partial X^{s}}\right),$
@@
$\mathrm{div}_{g}\left(\bm{v}\right)$ $=$ $\mathbf{P}\left[\begin{array}[]{c}\sum_{k=1}^{3}\mathcal{G}_{k}V^{1k}\\ \sum_{k=1}^{3}\mathcal{G}_{k}V^{2k}\\ \sum_{k=1}^{3}\mathcal{G}_{k}V^{3k}\end{array}\right]=\mathbf{P}\left[\begin{array}[]{ccc}\mathcal{G}_{1}&&\\ &\mathcal{G}_{1}&\\ &&\mathcal{G}_{1}\end{array}\right]\left[\begin{array}[]{c}V^{11}\\ V^{21}\\ V^{31}\end{array}\right]+\cdots+\mathbf{P}\left[\begin{array}[]{ccc}\mathcal{G}_{3}&&\\ &\mathcal{G}_{3}&\\ &&\mathcal{G}_{3}\end{array}\right]\left[\begin{array}[]{c}V^{13}\\ V^{23}\\ V^{33}\end{array}\right]$ (77)
$=$ $\mathcal{H}_{1}\left[\begin{array}[]{c}V^{11}\\ V^{21}\\ V^{31}\end{array}\right]+\mathcal{H}_{2}\left[\begin{array}[]{c}V^{12}\\ V^{22}\\ V^{32}\end{array}\right]+\mathcal{H}_{3}\left[\begin{array}[]{c}V^{13}\\ V^{23}\\ V^{33}\end{array}\right]:=\mathcal{H}_{1}\bm{V}\mathbf{|}_{1}+\mathcal{H}_{2}\bm{V}\mathbf{|}_{2}+\mathcal{H}_{3}\bm{V}\mathbf{|}_{3}$ (87)
where $\mathcal{H}_{s}$ is defined as in (37) and $\bm{V}\mathbf{|}_{s}:=(V^{1s},V^{2s},V^{3s})^{\top}$ is the $s$th column of tensor field $\bm{V}$ for $s=1,2,3$.
Let $\bm{v}={\mathrm{grad}}_{g}\bm{u}\in\mathfrak{X}(M)\times\mathfrak{X}(M)$ be the gradient of a vector field which is a $(2,0)$ tensor field. Using the extrinsic formulation in (4.1), one obtains the equation:
$\bm{V}=\left(\bm{V}\mathbf{|}_{1},\bm{V}\mathbf{|}_{2},\bm{V}\mathbf{|}_{3}\right)=\left(\mathcal{H}_{1}\bm{U},\mathcal{H}_{2}\bm{U},\mathcal{H}_{3}\bm{U}\right)=\mathbf{P}\left(\overline{{\mathrm{grad}}}_{\mathbb{R}^{3}}\bm{U}\right)\mathbf{P},$
@@
$\bm{V}\mathbf{|}_{k}=\mathcal{H}_{k}\bm{U},\text{ \ \ for }k=1,2,3\text{.}$ (88)
Combining (87) and (88), one can obtain the extrinsic formulation for the Bochner Laplacian of a vector field, $\Delta_{B}\bm{u}=g^{ij}u_{;ij}^{k}\frac{\partial}{\partial\xi^{k}}=\mathrm{div}_{g}({\mathrm{grad}}_{g}{\bm{u})}$, as
$\Delta_{B}\bm{u}$ $=$ $\mathrm{div}_{g}({\mathrm{grad}}_{g}\bm{u})=\mathrm{div}_{g}\left(\bm{v}\right)=\mathcal{H}_{1}\bm{V}\mathbf{|}_{1}+\mathcal{H}_{2}\bm{V}\mathbf{|}_{2}+\mathcal{H}_{3}\bm{V}\mathbf{|}_{3}$
$=$ $\mathcal{H}_{1}\mathcal{H}_{1}\bm{U}+\mathcal{H}_{2}\mathcal{H}_{2}\bm{U}+\mathcal{H}_{3}\mathcal{H}_{3}\bm{U}:=\bar{\Delta}_{B}\bm{U}.$